OpenStudy (anonymous):

Find the quotient. Write your answer in standard form. 2+i/2-i

OpenStudy (anonymous):

\[2+i/2-i =(2+i/2-i) * (2+i/2+i) = (4+4i-1/5)= (3/5) + (4/5) i\]